Two mirrors are inclined at an angle of 45degrees and a point object is placed between the two mirrors.the number of images formed is
1
Expert's answer
2020-12-21T11:01:07-0500

The number of images for a configuration of two mirrors inclined at angle θ is


"m=\frac{360°}{\theta}-1=\frac{360°}{45°}-1=7."